how do i simplify this equation or make the x the subject
6-(3-x)=21-7/2x
the 7/2 is a fraction

Answer by Alan3354(69443)

how do i simplify this equation or make the x the subject
6-(3-x)=21-7/2x
------------
I would say, "Solve for x."
6-(3-x)=21-7/2x
Multiply by 2 to eliminate fractions.
12 - 2(3-x) = 42 - 7x
12 -6 + 2x = 42 - 7x
6 + 2x = 42 - 7x
9x = 36
x = 4